# Ensemble Learning for Large Language Models in Text and Code Generation: A Survey

### v1

@@ 4.3 Results Aggregation Using Output Ensemble Models
   MoA-based open-source ensemble model with a deeper layer outperforms the single state-of-the-art closed-source LLM in instruction-following tasks.
4.4 Cost Effective Text Generation Using Routing Models
@@
This approach uses closed-source LLMs to achieve a cost-efficient performance that is 40% cheaper than that of a stronger LLM. It does this by modeling mixture-of-thought (MoT) representations.
A cascade-based language reasoning method was proposed in [9]. In this method, routing decisions are made based on the consistency of answers from a weaker LLM [105]. As explained in Section 3.2, an LLM cascade involves sequentially calling LLM APIs with prompts using a router, as illustrated in Figure 10.
To check for consistency, answers are first sampled from the weaker LLM using MoT representations. Each prompt includes CoT and PoT prompts, with eight examples for each type. This setup allows for 8-shot in-context learning. The consistency of the weaker LLM’s answers is assessed through voting or verification. The router uses this consistency score to decide whether to accept or reject the LLM’s answer. We evaluate both task accuracy and cost efficiency to understand their relationship.
A subjective reasoning test compared the performance of cascade-based models with stronger LLMs [9]. Figure 15 shows the cost-effective performance curves for these models. In this experiment, GPT models using CoT or PoT prompts and cascade-based models were compared. The results indicate that the cascade-based MoT model performed similarly to the GPT-4 model with CoT prompts, but at half the cost. Since coding requires reasoning skills, this cascade-based routing method could also generate optimal code at a lower cost.
The LLM routing approach, which combines weaker and stronger LLMs, can be applied to other text generation tasks, such as instruction-following [58, 57, 59], code generation [24], and MMLU [61]. For instruction-following tasks, routing-based methods have been developed to select the best output from various LLMs using reward models [58] or classifiers [57, 59]. In code generation and MMLU tasks, routing-based cost-effective methods have been benchmarked [24, 61]. The subjective evaluation results demonstrate the efficiency of this routing approach, achieving near-optimal performance at a low cost.
4.5 Comparisons Among These Four Approaches
The MoE approach with transformer models resembles traditional RNN-based language models. It maintains the input-to-expert routing step, requiring only minor changes to the existing language generation architecture for offline training [21]. Output ensemble methods, such as post-ranking fusion models or MoA models, expand a single LLM response into multiple responses [53, 23]. This allows for better capture of diversified long-range dependencies and improved generalization compared to conventional single LLM approaches. While training-based text generation from transformers can be complex [53], prompting LLMs for responses is simpler [23]. However, the computational time for LLMs can be slow due to the feed-forward multiLLM generation process with CoT prompting. Table 4 summarizes these recently proposed text and code generation approaches using LLM ensemble techniques. The model structure and advantage columns in the table also summarize different research directions. We analyze these approaches from three aspects based on their modeling traits and the relationship between input and output features. Further discussions on these approaches will be found in the ”Discussion” section.
@@ 4.5 Comparisons Among These Four Approaches
LLM Ensemble Method Application Model Structure Advantage Generated Features
A. MoE Code generation MoE transformer [21] Parameter efficient open-source LLM Next code token
Instruction-following MoE transformer [64] Next text token
Reasoning Switch transformer [6] Next text token
B. Output ensemble Instruction-following Ranker and fuser transformers [53] Extension flexible component Enhanced text tokens
C. Output ensemble with MoA Instruction-following Feed-forward structure [23] Training-free agentic workflow Synthesized response texts
D. Routing Reasoning LLM cascade + Decision maker [9] Cost effective Responded texts + Score
QA reasoning MoA + RF [55] Aggregated response texts
Instruction-following LLM routing + RL [58] Responded texts + Score
Instruction-following LLM routing + kNN [57] Responded texts + Score
Instruction-following LLM routing + transformer [59] Responded texts + Score
Code generation LLM routing + kNN or MLP [24] Responded code + Score
MMLU LLM routing + Light LLM [61] Responded texts + Score
   Model Structure and Advantage also summarize various research directions.
@@ 5 Discussion
5.1 Input and Target Features
In LLM ensembling for language generation, the input features depend on the model being used. Typically, language token embeddings serve as input features for transformer architectures [21, 53, 64, 6]. In a training-free model-level approach, we can only customize prompt engineering and agent structuring. Therefore, the input features for MoA and LLM cascade models usually consist of text prompts enhanced by prompt augmentation techniques [23, 55, 9].
In contrast, LLM routing models generally use text prompts or code snippets without prompt ensembling [58, 57, 59, 24, 61]. This suggests that routers should work with the original input to better classify which LLM to use.
Various output features for language generation are listed in Table 4. These features indicate that an effective combination of metrics is necessary, depending on the task and model. Section 2 explains that LLMs excel at capturing long-range dependencies in language features, which relate to human perception. Some LLM ensemble approaches consider this when evaluating their output text features, hinting at the potential for evaluating output code features as well. Human-perceptual correlations in text can be measured using metrics like BERTScore or BARTScore, comparing generated text to ground truth text [53, 57]. Similarly, LLM ensemble models can be evaluated in code generation by using CodeBERTScore [106] to compare generated code snippets to the ground truth.
In language generation with LLM cascade and routing models, the predicted score plays a crucial role in determining the final response [9, 58, 57, 59, 24, 61]. The quality of the final response from a routed LLM indicates the performance of this prediction.
@@ 5.2 Model Structures and Model Training
Table 4 shows that different model structures are used in these four approaches to achieve various goals. MoEs are employed for gating and expert functions, allowing for input-to-expert routing [21, 64, 6]. In contrast, post-ranking fusion transformers are used to represent input and output text features, enabling multioutput ranking and fusion [53].
In routing-based models, routers help invoke multiple LLMs sequentially [9] or route to the best single LLM [58, 57, 59, 24, 61]. They aim to produce cost-efficient outputs.
We recommend specific training or scoring options for each approach. The depth of the architecture, or the number of layers, is crucial for feed-forward ensemble models. In MoA-based instruction-following [23], the results in Table 3 indicate that increasing the number of layers improves the quality of synthetic responses due to more auxiliary information.
In other MoE studies [21, 64, 6], researchers tuned the number of transformer layers to minimize the cross-entropy loss between targets and outputs in training sets. The findings suggest that using more than 12 layers can enhance prediction accuracy. However, the optimal depth for MoE layers is typically not as deep as that used in switch-transformer-based text generation. This is reasonable, as $n$ the number of experts to be routed in switch transformer is $1$ while $n>1$ in MoE.
In routing-based approaches, various scoring strategies have been used. These include agreement scores to measure LLM consistency for reasoning [9], reward scores for instruction-following [58], and scoring methods using kNN, transformers, or LLMs for inference optimization [57, 59, 24, 61]. Given the high computational cost of training MoEs and other transformer models, one option is to skip the training process altogether by using a model-level agent ensemble [23, 55].
@@ 5.3 A Comparison Between Text and Code Generations Both Using LLM Ensemble Models
We discuss how LLM ensemble architectures are used in text generation and how they differ from those used in code generation. Understanding these differences can help us adapt text generation models for code generation tasks.
Transformers are the leading models for text generation. In this approach, a transformer is trained to convert an input sequence from an encoder (which processes natural language tokens) into an output sequence from a decoder. The decoder links the encoder’s representations to higher-level linguistic texts during generation.
LLM ensembling is commonly used for text generation. However, similar model structures have also been proposed for code generation, as shown in Table 4. Notably, the MoE and routing-based approaches for code generation [21, 24] use the same model structure as those used for text generation.
One of the key differences between the two is the training data. Text generation uses natural language data, wheres code generation relies on programming language data to fine-tune the LLM’s specialization.
Another difference lies in the evaluation metrics. For LLM-ensemble-based text generation, we typically measure instruction-following accuracy, answer correctness, and linguistic quality. In contrast, code generation focuses on functional correctness, runtime, and memory usage. Therefore, the metrics for code generation must reflect the specific requirements for evaluating code snippets. For instance, the pass rate of unit tests is a relevant metric for code generation but is not used in text generation tasks. However, some automatic metrics from text generation, like BERTScore [100] and BLEU [107], can be adapted for code generation, leading to metrics like CodeBERTScore [106] and CodeBLEU [108].
@@ 6 Conclusion
This article reviews new language generation methods that use LLM ensemble learning techniques. Compared to traditional methods that rely on a single GPT-based LLM, ensemble models at the architecture and model levels (e.g., MoEs, MoAs, post-ranking fusion, and routing) offer better representation of diversity among LLMs. They also combine output features more effectively, leading to improved naturalness, similarity to target responses, and overall quality of generated text or code.
We have examined various implementations of LLM ensemble techniques for text and code generation in current literature. To provide a clearer overview, we chose the effective approaches that could be applied in real-world scenarios. We grouped these approaches into four LLM ensemble methods and analyzed each method and connect findings systematically.
The findings reveal four key advantages of using these methods: 1) MoE models are parameter-efficient and open-source. They can outperform larger competitive LLMs with many more parameters. This makes them strong candidates for the future development of LLMs. 2) The output ensemble model utilizes a post-ranking fusion framework, consisting of a stand-alone ranker and fuser. This structure allows for flexibility in various extensions and applications. For instance, a conversational text fuser can be adapted to a code fuser for code generation, or a code ranker can score a list of code snippets. 3) The open-source MoA models have a training-free agentic workflow, making them easy to integrate and deploy. They can be flexibly adapted to different tasks by testing various prompts. 4) Routing provides a cost-effective way to manage powerful but expensive LLMs like GPT-4 and Gemini. This technique can greatly improve scalability, especially when handling thousands of prompts.
Many real-world tasks, such as chatbot assistance, medical diagnosis, mathematical reasoning, and code optimization, involve text or code generation. Therefore, our findings are relevant for model selection across various domains.
Despite the successes of LLM ensemble learning methods in text and code generation discussed in this article, important issues still need to be addressed to fully utilize the intrinsic strengths of LLM ensemble methods. For example, current methods have struggled to model and predict code tokens effectively using architectural LLM ensemble models [32, 20]. Given the differences between coding and natural language, we may need ensemble structures specifically designed for programming language modeling and prediction. Additionally, LLM ensemble approaches have not yet actively considered the generation optimality of multimodal features, including image, video, and acoustic features. We believe a promising direction for future research is to apply LLM ensemble models in a multimodal way. This research direction includes the models with better programming language modeling capabilities, such as abstract syntax trees, to code generation tasks and applying the LLM ensemble methods to multimodal LLMs in training and agentic workflow.

### v2

@@ 4.2.5 Output Ensemble
  MoA-based open-source ensemble model [23] outperforms the commercial GPT-4o-mini in instruction-following and code generation tasks.
4.2.6 Routing
Routing optimizes language model deployment by directing queries to the most suitable model from a pool of options, maximizing performance while minimizing costs. Recent advances [60, 61] demonstrate remarkable efficiency gains across diverse tasks (Table 7).
Key Findings. Routoo [61] shows that routing between open-source models can achieve 75.87% accuracy on MMLU, surpassing GPT-3.5 at less than half the cost. When including closed-source models, it approaches the performance of GPT-4-Turbo (84. 9% vs. 86. 4%) at half the cost. The multilabel classifier (MLC) router [60] similarly outperforms single open-source models. RouterBench [24] validates this approach in various domains, including code generation.
Resource Requirements. Routing systems use lightweight predictors to estimate model performance without full inference. Routoo implements a performance predictor and cost-aware selector with minimal overhead, enabling substantial cost savings while dynamically balancing quality, cost, and latency constraints based on deployment requirements.
Trade-offs. While routing systems offer impressive efficiency gains, their effectiveness depends on the prediction quality and the diversity of available models. Unlike ensembles that combine strengths across models, routing commits to a single model’s response per query. Predictors may struggle with novel task distributions not in their training data. Despite these limitations, routing provides an immediate deployable solution for optimizing LLM deployments without ensemble complexity or model fusion demands, particularly valuable for applications with heterogeneous query types.
Table 7: Routing Performance on text generation
[…]
Model MMLU Acc. Cost (\$/1M tok)
LLaMA2-13B 54.8% 0.26
Mistral-7B 62.09% 0.2
MLC Router [60] 63.85%
GPT3.5 70.0% 1.5
Routoo (open-source) [61] 75.87% 0.6
Routoo (mix) [61] 84.9% 10.2
GPT4-Turbo 86.40% 20
  The routing approach can significantly outperform single open-source LLMs [60]. Routing between open-source LLMs outperforms GPT-3.5 at less than half the cost [61]. Mixing in GPT-4 achieves performance comparable to GPT-4-Turbo at half the cost [61].
4.2.7 Cascading
Cascading approaches deploy multiple LLMs in sequence, routing simpler queries to cost-effective models while reserving expensive models for complex tasks. Pioneered in FrugalGPT [62] and enhanced with mixtures of thought representations [9], cascades achieve near-optimal performance while significantly reducing computational costs.
Key Findings. FrugalGPT achieves 1.5% higher accuracy than GPT-4 while reducing costs by 80% on classification tasks (Table 8). For reasoning tasks, models employing mixture of thought (MoT) representations match GPT-4’s performance at approximately half the cost, using answer consistency from multiple reasoning paths as routing signals between weaker and stronger models (Figure 12).
Resource Requirements. Cascade systems require minimal overhead, primarily an efficient routing mechanism based on consistency checks [9] or learned selectors [62]. This approach enables selective deployment of powerful models in cost-sensitive production environments.
Trade-offs. While cascades optimize costs, they introduce latency through sequential invocation and potential redundancy during query escalation. Their effectiveness depends on router accuracy—poor routing decisions can increase costs or degrade performance. Despite these challenges, cascading approaches demonstrate versatility across classification [62] and reasoning tasks [9], suggesting broader applicability in cost-constrained environments. Since coding requires reasoning skills, this cascade-based routing method could also generate optimal code at a lower cost.
Table 8: cascading performance and cost on classification
[…]
Model Accuracy Cost (\$)
GPT-4 0.857 33.1
FrugalGPT 0.872 6.5
  Cascading-based FrugalGPT [62] reduces the cost by 80%, while outperforms GPT-4 by 1.5% in classification tasks on the HEADLINES dataset.
@@ 4.2.7 Cascading
4.3 Comparative Analysis
LLM ensemble approaches offer distinct advantages based on deployment constraints and optimization goals. Weight merging techniques provide training-free enhancements ideal for quick specialization without computational overhead but typically require architectural compatibility [29, 31, 32]. Knowledge fusion accommodates structurally different models but demands more resources for fine-tuning [20, 44]. Mixture-of-Experts architectures excel in parameter efficiency by activating only relevant parameters, offering robust performance with controlled computational costs, though they require specialized pretraining [21, 101] unlike output ensemble [23] and cascading approaches [62, 9].
The selection of ensemble method depends primarily on constraints: computational efficiency favors weight merging and routing; architectural flexibility benefits from knowledge fusion and output ensembles; production cost optimization is best served by cascading and routing approaches; and human preference alignment typically requires reward ensembles. As LLM capabilities evolve, hybrid approaches may emerge as the most practical path to democratizing advanced language capabilities across diverse deployment scenarios, combining strengths while mitigating individual weaknesses of each ensemble technique.
Table 9 summarizes these recently proposed text and code generation approaches using LLM ensemble techniques. The model structure and advantage columns in the table also summarize different research directions. In the following ”Discussion” section, we further analyze these approaches from three aspects based on their metrics, modeling traits and the comparison between text and code generations.
@@ 4.3 Comparative Analysis
Ensemble Approach Method Model Structure Advantage Application
Weight Merging Task Arithmetic [29] code Pretrained LLMs Training-free performance enhancement Math reasoning
TIES Merging [31] code Coding
TIES + DARE [32] code Coding, instruction-following
EO [43] code Evolve weights + merging Automated model development Math reasoning
Knowledge Fusion FuseLLM [20] code Fine-tuned LLMs Can use structurally different models Reasoning, Coding
FusionChat [44]
Mixture-of-Experts Mixtral [21] code Pretrained transformer Parameter efficient robust performance MMLU, Coding
OLMOE [101] code
Reward Ensemble RLHF + LoRA [51] Fine-tuned LLMs + RL Superior in human alignment Conversation
RLHF + LoRA [48]
Output Ensemble MoA [23] code Feed-forward LLMs Training-free agentic workflow Instruction-following, Coding
LLM Blender [53] code Ranker and fuser transformers Extension flexible component Instruction-following
Routing MLC Router [60] code Multi-label classifier Cost effective without complexity MMLU
Routoo [61] code 2 lightweight LLMs
Cascading FrugalGPT [62] code Sequence of LLMs Cost effective with broader applications Classification
MoT [9] code MoT + Sequence of LLMs Math reasoning
  Model Structure and Advantage can summarize various research directions.
@@ 5 Discussion
5.1 Metrics
Metrics vary by task and model type. As discussed in Section 2, LLMs excel at capturing long-range dependencies in language features that correlate with human perception. This capability informs evaluation approaches for both text and code generation across LLM ensemble methods. For text generation, human-perceptual correlations can be measured using metrics like BERTScore [109] or BARTScore [94], which compare generated output to ground truth text [53]. Similarly, code generation quality can be evaluated using CodeBERTScore [110], which applies the same underlying principles to code snippets. In cascade and routing models, predicted quality scores determine which LLM’s response is selected as final output [9, 62, 24, 61], making response quality a direct indicator of routing effectiveness.
@@ 5.2 Model Structures and Model Training
Table 9 shows that different model structures are used in these seven approaches to achieve various goals. MoE employs pretrained transformers for gating and expert functions, allowing for input-to-expert routing [21, 101]. Similarly, weight merging uses pretrained LLMs to interpolate/extrapolate weights, enabling training-free performance enhancement [29, 31, 32]. Building on those pretrained models, knowledge fusion fine-tune them while fusing structurally different models [20, 44]. In contrast, reward ensemble uses those fine-tuned LLMs as backbones and combine them with RL approach for improved human alignment [51, 48].
Given the high computational cost of tuning LLMs and other transformer models, one option is to skip the training process altogether by using a model-level output ensemble. As a training-free agentic workflow, output ensemble models structure LLMs in a feed-forward way, enhancing performance comparable to commercial LLMs [23]. Alternatively, transformers can still be trained for multioutput ranking and fusion [53]. Both cases deal with LLM endpoints only.
In routing-based models, routers help invoke multiple LLMs sequentially [62, 9] or route to the best single LLM [60, 61]. They aim to produce cost-efficient outputs and use various scoring strategies. These include agreement scores to measure LLM consistency for reasoning [9], reward scores for instruction-following [58], and scoring methods using classifier, transformers, or LLMs for inference optimization [60, 59, 61].
@@ 5.3 A Comparison Between Text and Code Generations Both Using LLM Ensemble Models
While LLM ensemble architectures like weight-merging, knowledge fusion, MoE, and MoA share model structures across text and code generation [32, 20, 44, 21, 101, 23] (Table 9), their optimal application and the role of individual experts within the ensemble diverge significantly. These differences stem primarily from the nature of the generation target and the consequent evaluation paradigms.
A core distinction lies in the characteristics of the output. Text generation grapples with natural language’s inherent ambiguity, stylistic variety, and broad knowledge domains. Ensembles in this space often aim to synthesize diverse reasoning capabilities, blend stylistic nuances, or enhance robustness in complex instruction following. In contrast, code generation targets formal, structured programming languages where syntactic and semantic correctness are paramount for execution. Thus, ensemble models for code might specialize in ensuring logical integrity, generating specific valid constructs, or even different programming paradigms, with a stronger emphasis on verifiable correctness over subjective stylistic qualities.
These differing objectives are directly reflected in evaluation methodologies, which in turn influence ensemble design. For text, metrics often focus on instruction-following accuracy (e.g., AlpacaEval win rates), answer correctness (e.g., MMLU accuracy), and linguistic quality (e.g., BERTScore [109]), potentially leading to aggregation strategies that enhance fluency or consensus. Code generation, however, prioritizes functional correctness (e.g., pass@k on benchmarks like HumanEval), runtime, and memory usage. This frequently drives ensemble designs towards generating a diverse set of candidate solutions to maximize the probability of at least one passing unit tests, or selecting the best verifiable output. Metrics adapted from text generation, such as CodeBERTScore [110] and CodeBLEU [111], evaluate the similarity of generated code against a reference solution, capturing aspects like semantic equivalence even without exact token matches. However, this assessment of closeness to an exemplar provides a different signal than the direct confirmation of executional correctness and functional behavior offered by definitive pass/fail unit tests. Recognizing these distinct pressures is crucial for effectively adapting ensemble techniques from one domain to the other.
@@ 6 Conclusion
This review highlights recent LLM ensemble learning techniques for language generation. Methods such as weight merging, MoE, MoA, and routing enhance LLM diversity and output combination, improving the quality of generated text and code compared to single-model approaches. We systematically analyzed representative implementations for text and code generation to provide a clear overview of their real-world applicability.
Key advantages identified include:
1) Weight merging boosts performance without additional training. 2) Knowledge fusion enables fine-tuning of structurally different models. 3) Parameter-efficient MoE models can outperform much larger LLMs. 4) Training-free MoA models offer easy integration and task adaptability. 5) Routing and cascading provide cost-effective scalability for powerful LLMs like GPT-4 and Gemini. These findings are relevant for model selection in diverse applications like chatbot assistance, medical diagnosis, mathematical reasoning, and code optimization.
Despite these successes, challenges remain. Current weight merging and knowledge fusion methods show limitations in coding performance, suggesting a need for ensemble structures specifically designed for programming languages.
Furthermore, a significant frontier is extending ensemble principles to multimodal LLMs. While this review focused on text and code, the core concept of leveraging diverse model strengths is highly applicable to vision-language models like BLIP-2, Flamingo, and LLaVA. Future research should actively investigate ensembling multimodal outputs (e.g., via ranking and fusion), intelligent routing, knowledge fusion/weight merging across multimodal architectures, and specialized multimodal MoEs. Advancing ensemble strategies for both specialized code generation and the evolving landscape of multimodal AI is crucial for unlocking the full potential of collaborative intelligence in a wider range of applications.